# Gaussian Mixture Models Based Augmentation Enhances GNN Generalization

### v2

@@ 5 Conclusion
$\mathcal{R}(\ell_{aug})-\mathcal{R}(\ell)$ $=\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell_{aug}(\mathcal{G}_{n},\theta)\right|-\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell(\mathcal{G}_{n},\theta)\right|\right]$
$\leq\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell_{aug}(\mathcal{G}_{n},\theta)-\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell(\mathcal{G}_{n},\theta)\right|\right]$
$=\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\left(\ell_{aug}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right)\right|\right]$
$\leq\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\epsilon_{n}\left(\ell_{aug}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right)\right|\right]$
$\leq\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\ell_{aug}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right|$
$=\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta)-\ell(\mathcal{G}_{n},\theta)\right]\right|$
$\leq\max_{n\in\{1,\ldots,N\}}L_{Lip}\mathbb{E}_{\lambda\sim\mathcal{P}}\left[\left\|\mathcal{G}_{n}^{\lambda}-\mathcal{G}_{n}\right\|\right].$
Appendix B Proof of Theorem 3.3
In this section, we present the detailed proof of Theorem 3.3, which allows us to e perform an in-depth theoretical analysis of our augmentation strategy through the lens of influence functions.
Theorem 3.3 Given a test graph $\mathcal{G}_{k}$ from the test set, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in Equation 3. The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{i,j}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)$, where $\mathcal{G}_{n}^{\lambda_{n,m}}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{test}$ is given by
$\frac{d\ell(\mathcal{G}_{k}^{test},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}=-\nabla_{\theta}\ell(\mathcal{G}_{k}^{test},\hat{\theta})H_{\hat{\theta}}^{-1}\nabla_{\theta}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta}),$
where $\hat{\theta}_{\epsilon_{n,m}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{n,m}^{\text{aug}}$ denotes the parameters that minimize the upweighted objective function $\mathcal{L}_{n,m}^{\text{aug}}$ and $H_{\hat{\theta}}=\nabla_{\theta}^{2}\mathcal{L}(\hat{\theta})$ is the Hessian Matrix of the loss w.r.t the model parameters.
@@
Lets $\mathcal{G}_{n}^{\lambda_{n,m}}$ be an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter. The parameters $\hat{\theta}$ and $\hat{\theta}$ and $\hat{\theta}_{\epsilon_{n,m}}$ the parameters that minimize the empirical risk on the train set, i.e.,
$\hat{\theta}$ $=\arg\min_{\theta}\mathcal{L}.$
@@
$\hat{\theta}_{\epsilon_{n,m}}$ $=\arg\min_{\theta}\mathcal{L}_{n,m}^{\text{aug}}=\arg\min_{\theta}\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta).$
Therefore, we examine its firstorder optimality conditions,
@@
00 0 $=\nabla_{\hat{\theta}_{\epsilon_{n,m}}}\left(\mathcal{L}+\epsilon_{n,m}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\theta)\right).$ (7)
Using Taylor Expansion, we now develop the Equation 7. We have $\lim_{\epsilon_{n,m}\rightarrow 0}\hat{\theta}_{\epsilon_{n,m}}=\hat{\theta}$, thus,
$0\simeq\left[\nabla_{\hat{\theta}}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta})\right]+\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}^{2}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta})\right]\left(\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}\right).$
@@
$\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}^{2}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta})\right]^{-1}\left[\nabla_{\hat{\theta}}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta})\right].$
@@
$\frac{\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}}{\epsilon_{n,m}}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})\right]^{-1}\nabla_{\hat{\theta}}\ell(\mathcal{G}_{n}^{\lambda_{n,m}},\hat{\theta}).$

### v3

@@ Acknowledgment
$\mathcal{R}(\ell_{\text{aug}})-\mathcal{R}(\ell)$ $=\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell_{\text{aug}}(\mathcal{G}_{n},\theta)\right|-\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell(\mathcal{G}_{n},\theta)\right|\right]$
$\leq\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell_{\text{aug}}(\mathcal{G}_{n},\theta)-\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\ell(\mathcal{G}_{n},\theta)\right|\right]$
$=\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{n=1}^{N}\epsilon_{n}\left(\ell_{\text{aug}}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right)\right|\right]$
$\leq\mathbb{E}_{\epsilon_{n}\sim P_{\epsilon}}\left[\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\epsilon_{n}\left(\ell_{\text{aug}}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right)\right|\right]$
$\leq\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\ell_{\text{aug}}(\mathcal{G}_{n},\theta)-\ell(\mathcal{G}_{n},\theta)\right|$
$=\sup_{\theta\in\Theta}\frac{1}{N}\sum_{n=1}^{N}\left|\mathbb{E}_{\mathcal{G}_{n}^{\lambda}\sim A_{\lambda}}\left[\ell(\mathcal{G}_{n}^{\lambda},\theta)-\ell(\mathcal{G}_{n},\theta)\right]\right|$
$\leq\max_{n\in\{1,\ldots,N\}}\text{L}_{\text{Lip}}\mathbb{E}_{\mathcal{G}_{n}^{\lambda}\sim A_{\lambda}}\left[\left\|\mathcal{G}_{n}^{\lambda}-\mathcal{G}_{n}\right\|\right].$
Appendix B Proof of Proposition 3.2
In this section, we provide a detailed proof of the proposition 3.2, with the aim of providing sufficient theoretical conditions on the augmentation strategy $Q_{\lambda}$.
Proposition 3.2 Let $\delta_{\mathcal{D}}$ denote the discrete distribution of the training graph representations. Suppose we sample new augmented graph representations from a distribution $Q_{\lambda}$ defined on the support of $\delta_{\mathcal{D}}$. Then, the following inequality holds,
$\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]\leq\sqrt{2}\cdot\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{\mathbf{h}}\sim Q_{\lambda}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\left(\sqrt{\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})}+\sqrt{2}\right).$
where $\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})$ is is the Kullback-Leibler divergence from $\delta_{\mathcal{D}}$ to $Q_{\lambda}$.
@@
We have,
$\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]$ $=\int_{h\in\mathbb{R}^{d}}\int_{\widetilde{h}\in\mathbb{R}^{d}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\delta_{\mathcal{D}}(\mathbf{h})Q_{\lambda}(\widetilde{\mathbf{h}})\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}},$
where $\delta_{\mathcal{D}}:\mathbf{h}\mapsto\frac{1}{N}\sum_{n=1}^{N}\delta_{\mathbf{h}_{\mathcal{G}_{n}}}(\mathbf{h})$, and $\delta$ is the Dirac distribution.
$\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{\mathbf{h}}\sim Q_{\lambda}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]$ $=\int_{h\in\mathbb{R}^{d}}\int_{\widetilde{h}\in\mathbb{R}^{d}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\delta_{\mathcal{D}}(\mathbf{h})Q_{\lambda}(\widetilde{\mathbf{h}})\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}$
$\leq C\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\delta_{\mathcal{D}}(\mathbf{h})Q_{\lambda}(\widetilde{\mathbf{h}})\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}},$
where
$C=\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\\ h\neq\widetilde{h}\end{subarray}}\frac{\|\mathbf{h}-\widetilde{\mathbf{h}}\|\delta_{\mathcal{D}}(\mathbf{h})Q_{\lambda}(\widetilde{\mathbf{h}})}{\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|}.$
$\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}},\widetilde{h}\sim\mathcal{Q}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]$ $\leq C\int_{h\in\mathbb{R}^{d}}\int_{\widetilde{h}\in\mathbb{R}^{d}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}$
$\leq C\int_{h\in\mathbb{R}^{d}}\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}\\ h=\widetilde{h}\end{subarray}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}+C\int_{h\in\mathbb{R}^{d}}\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}\\ h\neq\widetilde{h}\end{subarray}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}.$
We first get an upperbound for the first term in the right side of the inequality,
[…]
$=\int_{h\in\mathbb{R}^{d}}\left(\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}h=\widetilde{h}\end{subarray}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(h)\right|\mathop{d\widetilde{\mathbf{h}}}\right)\mathop{d\mathbf{h}}$
$=\int_{h\in\mathbb{R}^{d}}\left(\int_{\widetilde{h}\in\mathbb{R}^{d}}\delta_{h}(\widetilde{h})\mathop{d\widetilde{\mathbf{h}}}\right)\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(h)\right|\mathop{d\mathbf{h}}$
$=\int_{h\in\mathbb{R}^{d}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(h)\right|\mathop{d\mathbf{h}}$
[…]
The term $\int_{h\in\mathbb{R}^{d}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q(h)\right|\mathop{d\mathbf{h}}$ corresponds to the Total Variation Distance (TD) $d_{TV}$ between $\delta_{\mathcal{D}}$ and $Q_{\lambda}$ (Yao & Liu, 2024), i.e.
[…]
Using Pinsker’s Inequality (Csiszár & Körner, 2011), we have,
$\int_{h\in\mathbb{R}^{d}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q(h)\right|\mathop{d\mathbf{h}}\leq\sqrt{2\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})}$
For the second term in the upperbound , we have,
$\int_{h\in\mathbb{R}^{d}}\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}\\ h\neq\widetilde{h}\end{subarray}}\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\delta_{\mathcal{D}}(\mathbf{h})\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}$ $=\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}\\ h\neq\widetilde{h}\end{subarray}}\left|\frac{1}{N}\sum_{n=N}^{N}\delta_{h_{\mathcal{G}_{n}}}(h)-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\left(\frac{1}{N}\sum_{n=N}^{N}\delta_{h_{\mathcal{G}_{n}}}(h)\right)\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}$
$=\frac{1}{N}\sum_{n=1}^{N}\int_{\begin{subarray}{c}\widetilde{h}\in\mathbb{R}^{d}\\ \widetilde{h}\neq h_{\mathcal{G}_{n}}\end{subarray}}\left|\frac{1}{N}-Q_{\lambda}(\widetilde{\mathbf{h}})\right|\mathop{d\mathbf{h}}\mathop{d\widetilde{\mathbf{h}}}\leq 2,$
because distributions are bounded between 0 and 1
Let now upperbound the constant $C$. We have $\forall a,b\in\mathbb{R}^{+},~{}ab\leq\frac{1}{2}(a-b)^{2}$. Therefore,
$C$ $=\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\\ h\neq\widetilde{h}\end{subarray}}\frac{\|\mathbf{h}-\widetilde{\mathbf{h}}\|\delta_{\mathcal{D}}(\mathbf{h})Q_{\lambda}(\widetilde{\mathbf{h}})}{\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|}$
$\leq\frac{1}{2}\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\\ h\neq\widetilde{h}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\left|\delta_{\mathcal{D}}(\mathbf{h})-Q_{\lambda}(\widetilde{\mathbf{h}})\right|$
$\leq\frac{1}{2}\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\\ h\neq\widetilde{h}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|2,~{}~{}\text{\qquad because distributions are bounded between 0 and 1,}$,
$=\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|.$
Therefore,
$\mathbb{E}_{\mathbf{h}\sim\delta_{\mathcal{D}}}\mathbb{E}_{\widetilde{h}\sim\mathcal{Q}}\left[\|\mathbf{h}-\widetilde{\mathbf{h}}\|\right]\leq\sqrt{2}\sup_{\begin{subarray}{c}\mathbf{h}\sim\delta_{\mathcal{D}}\\ \widetilde{h}\sim\mathcal{Q}\end{subarray}}\|\mathbf{h}-\widetilde{\mathbf{h}}\|\left(\sqrt{\mathcal{KL}(\delta_{\mathcal{D}}\parallel Q_{\lambda})}+\sqrt{2}\right).$
Appendix C Proof of Theorem 3.4
In this section, we present the detailed proof of Theorem 3.4, which allows us to perform an in-depth theoretical analysis of our augmentation strategy through the lens of influence functions.
@@
Theorem 3.4 Given a test graph $\mathcal{G}_{k}$ from the test set, let $\hat{\theta}=\arg\min_{\theta}\mathcal{L}$ be the GNN parameters that minimize the objective function in (1). The impact of upweighting the objective function $\mathcal{L}$ to $\mathcal{L}_{n,m}^{\text{aug}}=\mathcal{L}+\epsilon_{n,m}\ell(\widetilde{\mathcal{G}}_{n}^{m},\theta)$, where $\widetilde{\mathcal{G}}_{n}^{m}$ is an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter, on the model performance on the test graph $\mathcal{G}_{k}^{test}$ is given by
@@
$\frac{d\ell(\mathcal{G}_{k}^{\text{test}},\hat{\theta}_{\epsilon_{n,m}})}{d\epsilon_{n,m}}=-\nabla_{\theta}\ell(\mathcal{G}_{k}^{test},\hat{\theta})\mathbf{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}),$
where $\hat{\theta}_{\epsilon_{n,m}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{n,m}^{\text{aug}}$ denotes the parameters that minimize the upweighted objective function $\mathcal{L}_{n,m}^{\text{aug}}$ and $\mathbf{H}_{\hat{\theta}}=\nabla_{\theta}^{2}\mathcal{L}(\hat{\theta})$ is the Hessian Matrix of the loss w.r.t. the model parameters.
Proof.
Let $\widetilde{\mathcal{G}}_{n}^{m}$ be an augmented graph candidate of the training graph $\mathcal{G}_{n}$ and $\epsilon_{n,m}$ is a sufficiently small perturbation parameter. The parameters $\hat{\theta}$ and $\hat{\theta}_{\epsilon_{n,m}}$ the parameters that minimize the empirical risk on the train set, i.e.,
$\hat{\theta}$ $=\arg\min_{\theta}\mathcal{L},$
$\hat{\theta}_{\epsilon_{n,m}}$ $=\arg\min_{\theta}\mathcal{L}_{n,m}^{\text{aug}}=\arg\min_{\theta}\mathcal{L}+\epsilon_{n,m}\ell(\widetilde{\mathcal{G}}_{n}^{m},\theta).$
Therefore, we examine its first-order optimality conditions,
@@
00 0 $=\nabla_{\hat{\theta}_{\epsilon_{n,m}}}\left(\mathcal{L}+\epsilon_{n,m}\ell(\widetilde{\mathcal{G}}_{n}^{m},\theta)\right).$ (3)
Using Taylor Expansion, we now develop Eq. (3). We have $\lim_{\epsilon_{n,m}\rightarrow 0}\hat{\theta}_{\epsilon_{n,m}}=\hat{\theta}$, thus,
$0\simeq\left[\nabla_{\hat{\theta}}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta})\right]+\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}^{2}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta})\right]\left(\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}\right).$
@@
$\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}^{2}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta})\right]^{-1}\left[\nabla_{\hat{\theta}}\mathcal{L}(\hat{\theta})+\epsilon_{n,m}\nabla_{\hat{\theta}}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta})\right].$
@@
$\frac{\hat{\theta}_{\epsilon_{n,m}}-\hat{\theta}}{\epsilon_{n,m}}=-\left[\nabla_{\hat{\theta}}^{2}\mathcal{L}(\hat{\theta})\right]^{-1}\nabla_{\hat{\theta}}\ell(\widetilde{\mathcal{G}}_{n}^{m},\hat{\theta}).$